OpenStudy (anonymous):

Solve. 2/3(p+3)=1/3-P

OpenStudy (anonymous):

get rid of fraction! multiply both sides by 3

OpenStudy (anonymous):

\[\frac{2}{3}(p+3)=\frac{1}{3}-p\] \[3\left(\frac{2}{3}(p+3)\right)=3\times \left(\frac{1}{3}-p\right)\] \[2(p+3)=1-3p\]

OpenStudy (anonymous):

then proceed as usual \[2p+6=1-3p\]\[5p+6=1\] etc

OpenStudy (anonymous):

yay i was doing the process right :)

OpenStudy (anonymous):

let's see if i get it right

OpenStudy (anonymous):

Do you have to add 5p to each sides? @satellite73

OpenStudy (anonymous):

typo there after \[2p+6=1-3p\] add \(3p\) to both sides gives \[5p+6=1\] now subtact \(6\) from both sides

OpenStudy (anonymous):

you want all the \(p\) terms on one side, all numbers on the other you should get \(5p=-5\) then divide by 5

OpenStudy (anonymous):

would it be \[\frac{ -5 }{ 5 }\]

OpenStudy (anonymous):

Never mind got it's -1

OpenStudy (anonymous):

Thanks